# Two-dimensional delta Bose gas in a weighted space

### v1

@@ Acknowledgment
$\Big{\|}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{2m+1}\mathcal{T}^{\prime}_{k/2}(\tau_{k/2})\Big{\|}_{\mathrm{op}}\leq c^{m}\,m^{3}\,e^{\lambda t}\Big{(}c_{1}^{m}+c_{0}^{2}\Big{(}c_{1}+\frac{c_{0}^{2}}{\log(\lambda-c_{0}-1)}\Big{)}^{m-1}\Big{)}\ .$ (2.13)
Given Lemma 2.1, our task is reduced to bounding the norm $\|\cdot\|_{2,a\to 2,a}$ of the operators in (2.5). First, consider the multiplicative operator $\mathcal{E}:L^{2}(\mathbb{R}^{2})\to L^{2}_{a}(\mathbb{R}^{2})$, $\psi\mapsto e^{a|\cdot|_{1}}\psi$ and express the weights in the norm as conjugations:
$\big{\|}\mathcal{P}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $=\big{\|}\mathcal{E}^{2}\otimes\mathcal{E}^{\otimes[\![n]\!]\setminus\alpha}\cdot\mathcal{P}_{\alpha}(t)\cdot\mathcal{E}^{-\otimes[\![n]\!]}\big{\|}_{2\to 2}\ ,$ (2.14a)
$\big{\|}\mathcal{P}_{\alpha}(t)^{*}\big{\|}_{2,a\to 2,a}$ $=\big{\|}\mathcal{E}^{\otimes[\![n]\!]}\cdot\mathcal{P}_{\alpha}(t)^{*}\cdot\mathcal{E}^{-2}\otimes\mathcal{E}^{-\otimes[\![n]\!]\setminus\alpha}\big{\|}_{2\to 2}\ ,$ (2.14b)
$\big{\|}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\big{\|}_{2,a\to 2,a}$ $=\big{\|}\mathcal{E}^{2}\otimes\mathcal{E}^{\otimes[\![n]\!]\setminus\alpha}\cdot\mathcal{P}_{\alpha\alpha^{\prime}}(t)\cdot\mathcal{E}^{-2}\otimes\mathcal{E}^{-\otimes[\![n]\!]\setminus\alpha^{\prime}}\big{\|}_{2\to 2}\ ,$ (2.14c)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-\gamma t}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\Big{\|}_{2,a\to 2,a}$ $=\Big{\|}\mathcal{E}^{2}\otimes\mathcal{E}^{\otimes[\![n]\!]\setminus\alpha}\cdot\int_{0}^{\infty}\mathrm{d}t\,e^{-\gamma t}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\cdot\mathcal{E}^{-2}\otimes\mathcal{E}^{-\otimes[\![n]\!]\setminus\alpha^{\prime}}\Big{\|}_{2\to 2}\ ,$ (2.14d)
$\big{\|}\mathcal{J}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $=\big{\|}\mathcal{E}^{2}\otimes\mathcal{E}^{\otimes[\![n]\!]\setminus\alpha}\cdot\mathcal{J}_{\alpha}(t)\cdot\mathcal{E}^{-2}\otimes\mathcal{E}^{-\otimes[\![n]\!]\setminus\alpha}\big{\|}_{2\to 2}\ .$ (2.14e)
@@
$\mathsf{p}(t,x_{i}-y_{i})e^{a\,(|y_{i}|_{1}-|x_{i}|_{1})}\leq 2e^{2at}\mathsf{p}(2t,x_{i}-y_{i})\ ,\qquad x_{i},y_{i}\in\mathbb{R}^{2}\ .$ (2.15)
To see how, use the triangle inequality followed by Young’s inequality to get $|y_{i}|_{1}-|x_{i}|_{1}\leq|x_{i}-y_{i}|_{1}\leq 2at+|x-y|_{1}^{2}/(8at)\leq 2at+|x-y|_{2}^{2}/(4at)$, where we used the bound $|\cdot|_{1}^{2}\leq 2|\cdot|_{2}^{2}$ that holds on $\mathbb{R}^{2}$. This bounds the left-hand side of (2.15) by $\mathsf{p}(t,x_{i}-y_{i})\,e^{2at+|x_{i}-y_{i}|_{2}^{2}/(4t)}$. Further using the readily verified identity $\mathsf{p}(t,x_{i}-y_{i})e^{|x_{i}-y_{i}|_{2}^{2}/(4t)}=2\mathsf{p}(2t,x_{i}-y_{i})$ gives (2.15). Referring to (2.5a), we see that the conjugated operator in (2.14a) has kernel $e^{a|S_{\alpha}y|_{1}-a|x|_{1}}\mathcal{P}(t,S_{\alpha}y-x)$. Using (2.15) and the fact that the kernel is nonnegative, we bound the right-hand side of (2.14a) by $ce^{ct}\|\mathcal{P}_{\alpha}(2t)\|_{2\to 2}$. Hereafter, we write $c=c(\theta,n,a,\Phi)$ for a general, finite, positive, deterministic constant that may change from place to place but depend only on $\theta,n,a,\Phi$. Similar arguments apply to the operators in (2.14) and give, for $\alpha\neq\alpha^{\prime}$, the comparison bounds
$\big{\|}\mathcal{P}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}_{\alpha}(2t)\big{\|}_{2\to 2}\ ,$ (2.16a)
$\big{\|}\mathcal{P}_{\alpha}(t)^{*}\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}_{\alpha}(2t)^{*}\big{\|}_{2\to 2}\ ,$ (2.16b)
$\big{\|}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}_{\alpha\alpha^{\prime}}(2t)\big{\|}_{2\to 2}\ ,$ (2.16c)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-(c+2)t}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\Big{\|}_{2,a\to 2,a}$ $\leq c\,\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-2t}\mathcal{P}_{\alpha\alpha^{\prime}}(2t)\Big{\|}_{2\to 2}\ ,$ (2.16d)
$\big{\|}\mathcal{J}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\mathsf{j}(t)\,\|\mathsf{p}(t)\|_{2\to 2}\cdot\|\mathsf{p}(2t)\|_{2\to 2}^{n-2}\ .$ (2.16e)
We are now ready to prove Theorem 1.1(1). By [20, Lemmas  5.1, 8.1(a), 8.2, and 8.4] and the property that $\|\mathsf{p}(s)\|_{2\to 2}\leq 1$, the left-hand sides of (2.16) are bounded by $ce^{ct}t^{-1/2}$, $ce^{ct}t^{-1/2}$, $ce^{ct}t^{-1}$, $ce^{ct}$, and $ce^{ct}t^{-1}|\log(\frac{1}{2}\wedge t)|^{-2}$, respectively. Combining these bounds with Lemma 2.1 for a large enough $\lambda$ gives Theorem 1.1(1).
3. The prelimiting semigroup, proof of Theorem 1.1(2)
@@
$\mathbb{R}^{4}\times(\mathbb{R}^{2})^{[\![n]\!]\setminus\alpha}:=\mathbb{R}^{4}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}:=\big{\{}y=(y_{\mathrm{r}},y_{\mathrm{c}},(y_{i})_{i\in[\![n]\!]\setminus\alpha})\,\big{|}\,y_{\mathrm{r}},y_{\mathrm{c}},y_{i}\in\mathbb{R}^{2}\big{\}}\ ,$ (3.1)
$S^{\varepsilon}_{\alpha}:\mathbb{R}^{4}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}\to\mathbb{R}^{2[\![n]\!]},\qquad\big{(}S^{\varepsilon}_{\alpha}y\big{)}_{k}:=\begin{cases}y_{\mathrm{c}}+\varepsilon y_{\mathrm{r}}/2&\text{when }k=i\\ y_{\mathrm{c}}-\varepsilon y_{\mathrm{r}}/2&\text{when }k=j\\ y_{k}&\text{when }k\in[\![n]\!]\setminus\alpha\ .\end{cases}$ (3.2)
where we index the first two coordinates in (3.1) respectively by $\mathrm{r}$ and $\mathrm{c}$ for “relative” and “center of mass”. Set $\phi:=\sqrt{\Phi}$ and view $\Phi$ and $\phi$ as multiplicative operators acting on $L^{2}(\mathbb{R}^{2})=L^{2}(\{y_{\mathrm{r}}\})$. For $\alpha,\alpha^{\prime}\in\mathrm{Pair}[\![n]\!]$, define operators
$\mathcal{P}^{\varepsilon}_{\alpha}(t,y,x)$ $:=\phi(y_{\mathrm{r}})\,\mathcal{P}\big{(}t,S^{\varepsilon}_{\alpha}y-x\big{)}=:\big{(}\mathcal{P}^{\varepsilon}_{\alpha}\big{)}^{*}(t,x,y)\ ,$ (3.3a)
$\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t,y,y^{\prime})$ $:=\phi(y_{\mathrm{r}})\,\mathcal{P}^{\varepsilon}\big{(}t,S^{\varepsilon}_{\alpha}y-S^{\varepsilon}_{\alpha^{\prime}}y^{\prime}\big{)}\,\phi(y_{\mathrm{r}}^{\prime})\ ,$ (3.3b)
$\mathcal{J}^{\varepsilon}_{\alpha}(t)$ $:=\sum_{k=1}^{\infty}\beta_{\varepsilon}^{k+1}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\mathcal{P}^{\varepsilon}_{\alpha\alpha}(\tau_{1})\cdots\mathcal{P}^{\varepsilon}_{\alpha\alpha}(\tau_{k})\ ,$ (3.3c)
$\begin{split}\mathcal{R}^{\varepsilon}_{\vec{\alpha}}(t)&:=\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\mathcal{P}^{\varepsilon}_{\alpha_{1}}(\tau_{\frac{1}{2}})^{*}\prod_{k=1}^{|\vec{\alpha}|-1}\Big{(}\beta_{\varepsilon}\delta_{0}(\tau_{k})+\mathcal{J}^{\varepsilon}_{\alpha_{k}}(\tau_{k})\Big{)}\,\mathcal{P}^{\varepsilon}_{\alpha_{k}\alpha_{k+1}}(\tau_{k+\frac{1}{2}})\\ &\hskip 43.36464pt\cdot\Big{(}\beta_{\varepsilon}\delta_{0}(\tau_{|\vec{\alpha}|})+\mathcal{J}^{\varepsilon}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|})\Big{)}\,\mathcal{P}^{\varepsilon}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|+\frac{1}{2}})\ ,\end{split}$ (3.3d)
@@
$\mathsf{j}^{\varepsilon}(t):L^{2}(\mathbb{R}^{2})\to L^{2}(\mathbb{R}^{2})\ ,\quad\mathsf{j}^{\varepsilon}(t):=\sum_{\ell=1}^{\infty}\beta_{\varepsilon}^{\ell+1}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{\ell}\phi\,\mathsf{p}_{\varepsilon}(2\tau_{k})\,\phi.$ (3.4)
It is straightforward to check from (3.3c) that
$\mathcal{J}^{\varepsilon}_{\alpha}(t,y,y^{\prime})=\mathsf{j}^{\varepsilon}(t,y_{\mathrm{r}},y_{\mathrm{r}}^{\prime})\,\mathsf{p}(\tfrac{t}{2},y_{\mathrm{c}}-y_{\mathrm{c}}^{\prime})\,\prod_{i\in[\![n]\!]\setminus\alpha}\mathsf{p}(t,y_{i}-y^{\prime}_{i})\ .$ (3.5)
@@
$\mathcal{Q}^{\varepsilon}(t)=\mathcal{P}(t)+\sum_{\vec{\alpha}\in\mathrm{Dgm}[\![n]\!]}\mathcal{R}^{\varepsilon}_{\vec{\alpha}}(t)\ .$ (3.6)
To prove this, in (1.1), use Duhamel’s principle to write $\mathcal{Q}^{\varepsilon}(t)=\mathcal{P}(t)+\int_{0}^{t}\mathrm{d}s\,\mathcal{P}(t-s)\sum_{\alpha}\Phi^{\varepsilon}_{\alpha}\mathcal{Q}^{\varepsilon}(t)$ and iterate the equation to obtain
$\mathcal{Q}^{\varepsilon}(t)=\mathcal{P}(t)+\sum_{\vec{\eta}}\int_{\Sigma(t)}\mathrm{d}\tau\,\mathcal{P}^{\varepsilon}_{\eta_{1}}(\tau_{0})^{*}\prod_{k=2}^{|\vec{\eta}|}\mathcal{P}^{\varepsilon}_{\eta_{k-1}\eta_{k}}(\tau_{k})\cdot\mathcal{P}^{\varepsilon}_{\eta_{|\vec{\eta}|}}(\tau_{|\vec{\eta}|})\ ,$ (3.7)
where the sum runs over $\vec{\eta}\in\cup_{m=1}^{\infty}(\mathrm{Pair}[\![n]\!])^{m}$. Rewrite $\vec{\eta}$ as $(\alpha_{1}^{k_{1}},\alpha_{2}^{k_{2}},\ldots)$, where $\alpha^{k}:=(\alpha,\ldots,\alpha)\in(\mathrm{Pair}[\![n]\!])^{k}$, and $\alpha_{1}\neq\alpha_{2}$, $\alpha_{2}\neq\alpha_{3}$, …. Accordingly, the sum in (3.7) is rewritten as sums over $\vec{\alpha}\in\mathrm{Dgm}[\![n]\!]$ and over $k_{1},\ldots,k_{|\vec{\alpha}|}\in\mathbb{Z}_{>0}$. Carrying out the latter sum gives (3.6).
@@
$\|f\|_{L^{2}_{a}(\Omega)}^{2}:=\begin{cases}\int_{\Omega}\mathrm{d}x\,\big{|}f(x)e^{a|x|_{1}}\big{|}^{2}&\Omega=\mathbb{R}^{2[\![n]\!]}\ ,\\ \int_{\Omega}\mathrm{d}y\,\big{|}f(y)e^{a\sum_{\sigma=\pm 1}|y_{\mathrm{c}}+\frac{\varepsilon\sigma}{2}y_{\mathrm{r}}|_{1}+a\sum_{i\in[\![n]\!]\setminus\alpha}|y_{i}|_{1}}\big{|}^{2}&\Omega=\mathbb{R}^{4}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}\ .\end{cases}$ (3.8)
@@
We have slightly abused notation by using the same notation for the norms in (1.3) and in (3.8). This should not cause any confusion, since the latter will only be applied to $\varepsilon$-dependent operators. Define the operator norm $\|\cdot\|_{2,a\to 2,a}$ the same way as in (1.4) with the norm in (3.8) replacing that of (1.3).

### v2

@@ 1 Introduction
$\Big{\|}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{2m+1}\mathcal{T}^{\prime}_{k/2}(\tau_{k/2})\Big{\|}_{\mathrm{op}}\leq c^{m}\,m^{3}\,e^{\lambda t}\Big{(}c_{1}^{m}+c_{0}^{2}\Big{(}c_{1}+\frac{c_{0}^{2}}{\log(\lambda-c_{0}-1)}\Big{)}^{m-1}\Big{)}\ .$ (17)
Given Lemma 2.1, our task is reduced to bounding the norm $\|\cdot\|_{2,a\to 2,a}$ of the operators in (9). The Lemma then implies the convergence of the operator in equation (12) by choosing $\lambda$ large enough. First, consider the multiplicative isometric operator $\mathcal{E}:L^{2}_{a}(\mathbb{R}^{2})\to L^{2}(\mathbb{R}^{2})$, $\psi\mapsto e^{a|\cdot|_{1}}\psi$ and express the weights in the norm as conjugations:
$\big{\|}\mathcal{P}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $=\big{\|}\mathcal{E}^{2}\otimes\mathcal{E}^{\otimes[\![n]\!]\setminus\alpha}\cdot\mathcal{P}_{\alpha}(t)\cdot\mathcal{E}^{-\otimes[\![n]\!]}\big{\|}_{2\to 2}\ ,$ (18a)
$\big{\|}\mathcal{P}_{\alpha}(t)^{*}\big{\|}_{2,a\to 2,a}$ $=\big{\|}\mathcal{E}^{\otimes[\![n]\!]}\cdot\mathcal{P}_{\alpha}(t)^{*}\cdot\mathcal{E}^{-2}\otimes\mathcal{E}^{-\otimes[\![n]\!]\setminus\alpha}\big{\|}_{2\to 2}\ ,$ (18b)
$\big{\|}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\big{\|}_{2,a\to 2,a}$ $=\big{\|}\mathcal{E}^{2}\otimes\mathcal{E}^{\otimes[\![n]\!]\setminus\alpha}\cdot\mathcal{P}_{\alpha\alpha^{\prime}}(t)\cdot\mathcal{E}^{-2}\otimes\mathcal{E}^{-\otimes[\![n]\!]\setminus\alpha^{\prime}}\big{\|}_{2\to 2}\ ,$ (18c)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-\gamma t}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\Big{\|}_{2,a\to 2,a}$ $=\Big{\|}\mathcal{E}^{2}\otimes\mathcal{E}^{\otimes[\![n]\!]\setminus\alpha}\cdot\int_{0}^{\infty}\mathrm{d}t\,e^{-\gamma t}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\cdot\mathcal{E}^{-2}\otimes\mathcal{E}^{-\otimes[\![n]\!]\setminus\alpha^{\prime}}\Big{\|}_{2\to 2}\ ,$ (18d)
$\big{\|}\mathcal{J}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $=\big{\|}\mathcal{E}^{2}\otimes\mathcal{E}^{\otimes[\![n]\!]\setminus\alpha}\cdot\mathcal{J}_{\alpha}(t)\cdot\mathcal{E}^{-2}\otimes\mathcal{E}^{-\otimes[\![n]\!]\setminus\alpha}\big{\|}_{2\to 2}\ .$ (18e)
@@
$\mathsf{p}(t,x_{i}-y_{i})e^{a\,(|y_{i}|_{1}-|x_{i}|_{1})}\leq 2e^{2at}\mathsf{p}(2t,x_{i}-y_{i})\ ,\qquad x_{i},y_{i}\in\mathbb{R}^{2}\ .$ (19)
To see how, use the triangle inequality followed by Young’s inequality to get $|y_{i}|_{1}-|x_{i}|_{1}\leq|x_{i}-y_{i}|_{1}\leq 2at+|x-y|_{1}^{2}/(8at)\leq 2at+|x-y|_{2}^{2}/(4at)$, where we used the bound $|\cdot|_{1}^{2}\leq 2|\cdot|_{2}^{2}$ that holds on $\mathbb{R}^{2}$. This bounds the left-hand side of (19) by $\mathsf{p}(t,x_{i}-y_{i})\,e^{2at+|x_{i}-y_{i}|_{2}^{2}/(4t)}$. Further using the readily verified identity $\mathsf{p}(t,x_{i}-y_{i})e^{|x_{i}-y_{i}|_{2}^{2}/(4t)}=2\mathsf{p}(2t,x_{i}-y_{i})$ gives (19). Referring to (9a), we see that the conjugated operator in (18a) has kernel $e^{a|S_{\alpha}y|_{1}-a|x|_{1}}\mathcal{P}(t,S_{\alpha}y-x)$. Using (19) and the fact that the kernel is nonnegative, we bound the right-hand side of (18a) by $ce^{ct}\|\mathcal{P}_{\alpha}(2t)\|_{2\to 2}$. Hereafter, we write $c=c(\theta,n,a,\Phi)$ for a general, finite, positive, deterministic constant that may change from place to place but depends only on $\theta,n,a,\Phi$. Similar arguments apply to the operators in (18) and give, for $\alpha\neq\alpha^{\prime}$, the comparison bounds
$\big{\|}\mathcal{P}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}_{\alpha}(2t)\big{\|}_{2\to 2}\ ,$ (20a)
$\big{\|}\mathcal{P}_{\alpha}(t)^{*}\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}_{\alpha}(2t)^{*}\big{\|}_{2\to 2}\ ,$ (20b)
$\big{\|}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\big{\|}\mathcal{P}_{\alpha\alpha^{\prime}}(2t)\big{\|}_{2\to 2}\ ,$ (20c)
$\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-(c+2)t}\mathcal{P}_{\alpha\alpha^{\prime}}(t)\Big{\|}_{2,a\to 2,a}$ $\leq c\,\Big{\|}\int_{0}^{\infty}\mathrm{d}t\,e^{-2t}\mathcal{P}_{\alpha\alpha^{\prime}}(2t)\Big{\|}_{2\to 2}\ ,$ (20d)
$\big{\|}\mathcal{J}_{\alpha}(t)\big{\|}_{2,a\to 2,a}$ $\leq ce^{ct}\mathsf{j}(t)\,\|\mathsf{p}(t)\|_{2\to 2}\cdot\|\mathsf{p}(2t)\|_{2\to 2}^{n-2}\ .$ (20e)
We are now ready to prove Theorem 1.1(1). By [20, Lemmas  5.1, 8.1(a), 8.2, and 8.4] and the property that $\|\mathsf{p}(s)\|_{2\to 2}\leq 1$, the left-hand sides of (20) are bounded by $ce^{ct}t^{-1/2}$, $ce^{ct}t^{-1/2}$, $ce^{ct}t^{-1}$, $ce^{ct}$, and $ce^{ct}t^{-1}|\log(\frac{1}{2}\wedge t)|^{-2}$, respectively. Combining these bounds with Lemma 2.1 for a large enough $\lambda$ gives Theorem 1.1(1).
3 The prelimiting semigroup, proof of Theorem 1.1(2)
@@
$\mathbb{R}^{4}\times(\mathbb{R}^{2})^{[\![n]\!]\setminus\alpha}:=\mathbb{R}^{4}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}:=\big{\{}y=(y_{\mathrm{r}},y_{\mathrm{c}},(y_{i})_{i\in[\![n]\!]\setminus\alpha})\,\big{|}\,y_{\mathrm{r}},y_{\mathrm{c}},y_{i}\in\mathbb{R}^{2}\big{\}}\ ,$ (21)
$S^{\varepsilon}_{\alpha}:\mathbb{R}^{4}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}\to\mathbb{R}^{2[\![n]\!]},\qquad\big{(}S^{\varepsilon}_{\alpha}y\big{)}_{k}:=\begin{cases}y_{\mathrm{c}}+\varepsilon y_{\mathrm{r}}/2&\text{when }k=i\\ y_{\mathrm{c}}-\varepsilon y_{\mathrm{r}}/2&\text{when }k=j\\ y_{k}&\text{when }k\in[\![n]\!]\setminus\alpha\ .\end{cases}$ (22)
where we index the first two coordinates in (21) respectively by $\mathrm{r}$ and $\mathrm{c}$ for “relative” and “center of mass”. Set $\phi:=\sqrt{\Phi}$ and view $\Phi$ and $\phi$ as multiplicative operators acting on $L^{2}(\mathbb{R}^{2})$, where $\mathbb{R}^{2}$ represents the space of the $y_{\mathrm{r}}$ coordinate. For $\alpha,\alpha^{\prime}\in\mathrm{Pair}[\![n]\!]$, define operators
$\mathcal{P}^{\varepsilon}_{\alpha}(t,y,x)$ $:=\phi(y_{\mathrm{r}})\,\mathcal{P}\big{(}t,S^{\varepsilon}_{\alpha}y-x\big{)}=:\big{(}\mathcal{P}^{\varepsilon}_{\alpha}\big{)}^{*}(t,x,y)\ ,$ (23a)
$\mathcal{P}^{\varepsilon}_{\alpha\alpha^{\prime}}(t,y,y^{\prime})$ $:=\phi(y_{\mathrm{r}})\,\mathcal{P}\big{(}t,S^{\varepsilon}_{\alpha}y-S^{\varepsilon}_{\alpha^{\prime}}y^{\prime}\big{)}\,\phi(y_{\mathrm{r}}^{\prime})\ ,$ (23b)
$\mathcal{J}^{\varepsilon}_{\alpha}(t)$ $:=\sum_{k=1}^{\infty}\beta_{\varepsilon}^{k+1}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\mathcal{P}^{\varepsilon}_{\alpha\alpha}(\tau_{1})\cdots\mathcal{P}^{\varepsilon}_{\alpha\alpha}(\tau_{k})\ ,$ (23c)
$\begin{split}\mathcal{R}^{\varepsilon}_{\vec{\alpha}}(t)&:=\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\mathcal{P}^{\varepsilon}_{\alpha_{1}}(\tau_{\frac{1}{2}})^{*}\prod_{k=1}^{|\vec{\alpha}|-1}\Big{(}\beta_{\varepsilon}\delta_{0}(\tau_{k})+\mathcal{J}^{\varepsilon}_{\alpha_{k}}(\tau_{k})\Big{)}\,\mathcal{P}^{\varepsilon}_{\alpha_{k}\alpha_{k+1}}(\tau_{k+\frac{1}{2}})\\ &\hskip 43.36464pt\cdot\Big{(}\beta_{\varepsilon}\delta_{0}(\tau_{|\vec{\alpha}|})+\mathcal{J}^{\varepsilon}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|})\Big{)}\,\mathcal{P}^{\varepsilon}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|+\frac{1}{2}})\ ,\end{split}$ (23d)
@@
$\mathsf{j}^{\varepsilon}(t):L^{2}(\mathbb{R}^{2})\to L^{2}(\mathbb{R}^{2})\ ,\quad\mathsf{j}^{\varepsilon}(t):=\sum_{\ell=1}^{\infty}\beta_{\varepsilon}^{\ell+1}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{\ell}\phi\,\mathsf{p}_{\varepsilon}(2\tau_{k})\,\phi.$ (24)
It is straightforward to check from (23c) that
$\mathcal{J}^{\varepsilon}_{\alpha}(t,y,y^{\prime})=\mathsf{j}^{\varepsilon}(t,y_{\mathrm{r}},y_{\mathrm{r}}^{\prime})\,\mathsf{p}(\tfrac{t}{2},y_{\mathrm{c}}-y_{\mathrm{c}}^{\prime})\,\prod_{i\in[\![n]\!]\setminus\alpha}\mathsf{p}(t,y_{i}-y^{\prime}_{i})\ .$ (25)
@@
$\mathcal{Q}^{\varepsilon}(t)=\mathcal{P}(t)+\sum_{\vec{\alpha}\in\mathrm{Dgm}[\![n]\!]}\mathcal{R}^{\varepsilon}_{\vec{\alpha}}(t)\ .$ (26)
To prove this, in (1), use Duhamel’s principle to write $\mathcal{Q}^{\varepsilon}(t)=\mathcal{P}(t)+\int_{0}^{t}\mathrm{d}s\,\mathcal{P}(t-s)\sum_{\alpha}\Phi^{\varepsilon}_{\alpha}\mathcal{Q}^{\varepsilon}(t)$ and iterate the equation to obtain
$\mathcal{Q}^{\varepsilon}(t)=\mathcal{P}(t)+\sum_{\vec{\eta}}\int_{\Sigma(t)}\mathrm{d}\tau\,\mathcal{P}^{\varepsilon}_{\eta_{1}}(\tau_{0})^{*}\prod_{k=2}^{|\vec{\eta}|}\mathcal{P}^{\varepsilon}_{\eta_{k-1}\eta_{k}}(\tau_{k})\cdot\mathcal{P}^{\varepsilon}_{\eta_{|\vec{\eta}|}}(\tau_{|\vec{\eta}|})\ ,$ (27)
where the sum runs over $\vec{\eta}\in\cup_{m=1}^{\infty}(\mathrm{Pair}[\![n]\!])^{m}$. Rewrite $\vec{\eta}$ as $(\alpha_{1}^{k_{1}},\alpha_{2}^{k_{2}},\ldots)$, where $\alpha^{k}:=(\alpha,\ldots,\alpha)\in(\mathrm{Pair}[\![n]\!])^{k}$, and $\alpha_{1}\neq\alpha_{2}$, $\alpha_{2}\neq\alpha_{3}$, …. Accordingly, the sum in (27) is rewritten as sums over $\vec{\alpha}\in\mathrm{Dgm}[\![n]\!]$ and over $k_{1},\ldots,k_{|\vec{\alpha}|}\in\mathbb{Z}_{>0}$. Carrying out the latter sum gives (26).
@@
$\|f\|_{L^{2}_{a}(\Omega)}^{2}:=\begin{cases}\int_{\Omega}\mathrm{d}x\,\big{|}f(x)e^{a|x|_{1}}\big{|}^{2}&\Omega=\mathbb{R}^{2[\![n]\!]}\ ,\\ \int_{\Omega}\mathrm{d}y\,\big{|}f(y)e^{a\sum_{\sigma=\pm 1}|y_{\mathrm{c}}+\frac{\varepsilon\sigma}{2}y_{\mathrm{r}}|_{1}+a\sum_{i\in[\![n]\!]\setminus\alpha}|y_{i}|_{1}}\big{|}^{2}&\Omega=\mathbb{R}^{4}\times\mathbb{R}^{2[\![n]\!]\setminus\alpha}\ .\end{cases}$ (28)
@@
We have slightly abused notation by using the same notation for the norms in (3) and in (28). This should not cause any confusion, since the latter will only be applied to $\varepsilon$-dependent operators. Define the operator norm $\|\cdot\|_{2,a\to 2,a}$ the same way as in (4) with the norm in (28) replacing that of (3).